OpenStudy (anonymous):

Jose has $2.50 in nickels and dimes in his jar. If he has 11 more nickels than dimes, how many dimes does he have?

OpenStudy (anonymous):

Let d= the number of dimes; Let n= the number of nickels. Since there are 11 more nickels than dimes then the number of nickels is 11+d=n The value of the number of dimes + the value of the number of nickels is equal to $2.50. So we have this equation. d(.10)+n(.05)=2.50 Substitution d+11 for n we get d(.10)+(d+11)(.05)=2.50 Which gives us .1d+.55+.05d=2.50 Combining terms gives us .15d+.55=2.50 Subtracting .55 from both sides gives us .15d=1.95 Dividing both sides by .15 gives us d=13 dimes.